OpenStudy (anonymous):

There are 2 blue cards, 3 green cards, 1 yellow card, and 4 pink cards. What is the probability of drawing a green card, NOT replacing it, and then drawing another green card? A. 9/100 B. 3/50 C. 1/15

jimthompson5910 (jim_thompson5910):

P(green and green) = P(green)*P(green) P(green and green) = (3/10)*(2/9) P(green and green) = 6/90 P(green and green) = 1/15 So the answer is choice C